# Small Batch Size Training for Language Models: When Vanilla SGD Works, and Why Gradient Accumulation Is Wasteful

### v3

@@ Abstract
Figure 1: Small batch sizes are more robust to optimizer design and hyperparameter values. (a) Loss achieved by a transformer decoder-only language model with 30M active parameters trained on 600M tokens of FineWeb-Edu data using SGD, Adafactor, Adam, and Muon. At small batch sizes, all optimizers achieve similar loss; at large batch sizes, the gap between optimizers grows. (b) Loss achieved by GPT-2 (124M) trained on 2.5B tokens of FineWeb using AdamW when tuning the learning rate and $\beta_{1}$ hyperparameters for batch sizes 1 and 512. While both batch sizes achieve a similar lowest loss, the smaller batch size is much more robust to the hyperparameter values, reducing the need for hyperparameter tuning. For batch size 512, we use the default hyperparameters from Brown et al. [1] and the nanoGPT public repository [2]. For batch size 1, we turn off weight decay and rescale $\beta_{2}$ to preserve the token half-life, resulting in $\beta_{2}=0.9999$.
@@ 1 Introduction
Large batch sizes are widely believed to improve the stability of language model training [3, 4, 5, 6]. As a consequence, sophisticated optimizers that perform well in large-batch training are increasingly standard practice [7, 8, 9, 10, 11, 12]. In fact, it is common to simulate batch sizes even larger than the maximum batch size that fits into device memory through gradient accumulation [6, 5, 13, 14].
In small batch size pretraining experiments, one may observe loss spikes and heavy instability [15, 16, 17, 18]. While it is common to decrease the learning rate for smaller batch sizes, other hyperparameters, such as the decay rates for the first and second moments in Adam ($\beta_{1}$ and $\beta_{2}$), are often held fixed across batch sizes. We show that if instead of holding $\beta_{2}$ fixed, we hold the half-life of $\beta_{2}$ fixed (measured in number of tokens), stable training is possible all the way down to batch size one, as illustrated in Figure˜1(a).
When scaling hyperparameters in this way, we find that small batch sizes can confer computational advantages and greater robustness. In the small batch regime, we observe that the speed of convergence is less sensitive to optimizer hyperparameters like the learning rate, and momentum becomes less necessary, as we find in Figure˜1(b). In contrast, large batch sizes require large learning rates to compensate for taking fewer steps at a given compute budget. Large step sizes in turn require that the optimizer make predictions about the loss surface far away from the current iterate. We find that predicting such far-away parameter updates requires a sophisticated and carefully tuned optimizer. For very small batch sizes, we find that even vanilla stochastic gradient descent (SGD), with no momentum or weight decay, becomes competitive for training language models, in contrast to findings from recent work [19]. The ability to perform stable training without momentum can confer significant memory advantages, since we do not need to store the corresponding optimizer state, which can have a significant memory footprint. Based on our observations, we also revisit Adafactor, a memory-efficient variant of Adam, and show that it can be a compelling alternative to Adam in the small batch regime, enabling training of larger models in a memory-constrained setting.
@@
It is common in convex optimization to use higher batch sizes as the loss converges to a minimum [3, 20]. Intuitively, the higher the gradient to gradient noise magnitude ratio, the smaller the batch size we should take. For example, if the gradient noise magnitude is close to zero, then we obtain roughly the same gradient estimate with a small batch size as a large batch size but pay far fewer floating-point operations (FLOPs), so small batch sizes are efficient. Smaller batch sizes allow us to take more steps for a fixed FLOP budget. As we converge to a minimum, that same ratio instead often tends to zero since the gradient norm vanishes while the gradient noise may not, in which case we may want a larger batch size. We hypothesize that language model training lives in the far-from-convergence regime where small batch sizes are efficient. Following compute-optimal scaling laws, we would not train language models into the convergence regime since we could achieve better performance by instead training a larger model on less data [21, 22].
@@ 2 Background: Optimization for Language Models
Stochastic gradient descent (SGD) [23, 24]. SGD updates model parameters by computing gradients on mini-batches of data as follows: $\theta_{t+1}=\theta_{t}-\eta g_{t}$, where $\theta_{t}$ are the model parameters at step $t$, $\eta$ is the learning rate, and $g_{t}=\nabla_{\theta}\mathcal{L}(\theta_{t})$ is the gradient of the loss function with respect to $\theta_{t}$ averaged over $B$ samples. We refer to $B$ as the batch size and consider the extreme case of $B=1$ in our experiments.
@@
Adam [25]. Adam is an adaptive optimizer that maintains an exponential moving average (EMA) of the gradient and squared gradient, referred to as the first and second moments, and denoted as $m_{t}$ and $v_{t}$. The timescales of these moving averages are controlled by decay rates $\beta_{1}$ and $\beta_{2}$:
@@
Higher values of $\beta_{1}$ and $\beta_{2}$ place more weight on past gradients, meaning that the moments are averaged over longer timescales and evolve slowly. When training language models, values like $\beta_{1}=0.9$ and $\beta_{2}\in[0.95,0.98]$ are often used in practice [1, 26]. While most researchers and practitioners do not scale $\beta_{1}$ and $\beta_{2}$ with the batch size, we show in our work that scaling $\beta_{2}$ in particular is crucial for achieving good performance with Adam at a small batch size.
@@
Adafactor [27]. The second moment estimate $v_{t}$ used in Adam requires storing a number of floating point values equal to the parameter count of the model itself. Instead, Adafactor only stores the per-row and per-column sums of the moving average for the neural network weight matrices and estimates the per-parameter second moment using these sums. Adafactor also does not store a first moment estimate at all. Therefore, for a weight matrix with shape $d_{1}\times d_{2}$, the memory requirements for the moving averages are reduced from $2\times d_{1}\times d_{2}$ for Adam to $d_{1}+d_{2}$ for Adafactor. We demonstrate in Section˜4.5 that thanks to this sublinear memory cost, Adafactor achieves the best performance–memory trade-off.
Muon [11]. Another optimizer that was recently demonstrated to be competitive for language model training is MomentUm Orthogonalized by Newton-Schulz (Muon). Muon is specifically designed for $\geq 2$-dimensional parameter tensors, such as weight matrices in linear layers, where the Newton-Schulz iterative method is applied to the first moment estimate $m_{t}$ before using it to update the parameters. Muon is often used in conjunction with other optimizers like Adam. In our experiments using Muon, we only apply it to hidden layers and use Adam for input embedding and final layers.
@@ 3 Related Work
Batch size has long been studied in both theory and practice: its impact on the optimization path [28, 29, 30, 31], its interaction with different optimizers [32, 33, 34, 35], and its effect on the optimal learning rates and momentum hyperparameters [36, 37, 38, 39, 34, 40, 41, 42, 41, 43, 44]. Prior work has also explored related themes, such as the flatness argument favoring SGD [45, 46, 47, 48], SGD’s implicit biases [49, 50, 51, 52], and the comparison between stochastic and full-batch optimization [53, 54, 55, 56]. However, modern language model training is neither convex nor characterized by multiple epoch training like vision settings. In fact, we often train for a limited token budget and stop far before convergence, since otherwise scaling laws suggest that we could train a much better language model for the same compute budget by training a larger model on fewer tokens [22]. Hence, prior intuitions, such as the need for a larger batch size in the convergence phase, may not apply. For this reason, in the rest of this section, we cover works that focus on modern language model training specifically.
@@
Zhang et al. [41] and Shallue et al. [34] examine the critical batch size, which represents the threshold beyond which greater data parallelism would lead to diminishing returns, and tune all optimizer hyperparameters. They argue that any batch size below the critical batch size should perform equally well when training for the same number of tokens. Differently from these works, we empirically demonstrate that small batch sizes, including the extreme value of batch size one, perform on par with or even better than larger batch sizes, and we also show that small batches are more robust to optimizer and hyperparameter choices and enable SGD to perform competitively. We validate these findings on language models with modern design choices and up to 1.3 billion parameters.
@@
On the other hand, Vyas et al. [57] argue that gradient noise from small batch sizes does not yield similar performance advantages in the single epoch setting as it does in multiple-epoch training. Experimental results from Filatov et al. [58] also suggest that small batches do not work as well as large batches for language models trained on C4 [59], but notably, the authors do not tune the decay rates $\beta_{1}$ and $\beta_{2}$ of the Adam optimizer. In the same vein, Xiao [16] claim that both excessively small and large batch sizes underperform for language model training with Adam, leading to a U-shaped curve of the loss as a function of the batch size. We reproduce the exact experiments performed by Xiao [16] and show that their conclusions are an artifact of incomplete hyperparameter tuning; we find that simply using a slower decay rate of the second moment estimate $\beta_{2}$ makes small batch sizes competitive.
@@
Zhao et al. [19] train language models on C4 using different diagonal preconditioning optimizers, namely Adam, Lion, Adafactor with momentum, SGD with momentum, and signed SGD with momentum. They demonstrate that all these optimizers, aside from SGD which lags behind, achieve comparable optimal performance and robustness to hyperparameters. We challenge this conclusion in our work and argue that in the small batch size regime, SGD can perform on par with Adam.
@@
Shallue et al. [34], Kidambi et al. [40] and Zhang et al. [41] find that momentum is not necessary for SGD when the batch size is small, which aligns with our findings. However, Kidambi et al. [40] focus on small-scale vision experiments, Shallue et al. [34] and Kidambi et al. [40] do not compare the performance of SGD vs. Adam, and Zhang et al. [41] claim that SGD performs significantly worse than Adam. We hypothesize this observation could be a result of Zhang et al. [41] using 64 as the smallest batch size – we test batch sizes all the way down to 1. Zhang et al. [35] highlight that the benefits of more sophisticated optimizers diminish as batch size decreases, but their theoretical results assume a convex quadratic objective and their empirical results only cover a single language modeling experiment that is limited to a two-layer transformer.
@@
Porian et al. [42] and Zhang et al. [41] find that increasing the value of $\beta_{2}$ is important for small-batch training with Adam. However, their approach is based on discrete hyperparameter sweeps across $\beta_{2}$. In contrast, we introduce and validate a principled scaling rule that holds the second-moment half-life constant in terms of tokens. Busbridge et al. [60] propose a scaling rule for the decay coefficient of exponential moving averages similar to ours, but applied in the context of model weight averaging. Zhang et al. [44] show that there exists a threshold of $\beta_{2}^{*}$ such that when $\beta_{2}$ > $\beta_{2}^{*}$, Adam converges (this threshold decreases with batch size, i.e. the threshold is higher for smaller batch sizes). Chowdhery et al. [43] note that increasing $\beta_{2}$ during training leads to better performance for rare token embeddings. While this is an important insight, we provide a simpler batch-size-aware heuristic that applies even when using a constant $\beta_{2}$ throughout training.
@@ 4 Experimental Results
Recent works find that larger batch sizes and relatively sophisticated optimizers are necessary for stable language model training. For instance, Zhao et al. [19], Kunstner et al. [61], and Zhang et al. [62] observe that SGD has an extremely slow convergence speed for training language models in comparison to adaptive optimizers like Adam. Other works find that Adam performs poorly with small batch sizes [57, 16, 63, 64].
In this section, we revisit and challenge these beliefs by noting that such findings arise due to poor hyperparameter choices and specific experimental setups, for example, not adjusting the decay parameters $\beta_{1}$ and $\beta_{2}$ for Adam at small batch sizes, or using too large a batch size for SGD. We put these beliefs to the test by running a thorough grid search over optimizer hyperparameters and batch sizes for SGD, Adam, Adafactor, and Muon. As the cost of a grid search grows exponentially with the number of hyperparameters, we only perform exhaustive grid searches on a small transformer decoder-only model with 30 million active parameters following Liu et al. [65] and Xiao [16]. We use these results to propose heuristics for scaling hyperparameters across batch sizes, and we validate our findings at a larger scale by pretraining GPT-2 (124M) [66] and GPT-3 (1.3B) [1] and fine-tuning Gemma 3 (4B) [67]. In our experiments, we study batch sizes spanning more than four orders of magnitude: $\{1,4,16,64,256,1024,\text{ and }4096\}$.
@@ 4.1 Small Batch Sizes Render Sophisticated Optimizers Unnecessary
We train the 30M model on 600 million tokens from the FineWeb-Edu dataset [68], following Chinchilla scaling laws [22]. Our model adopts modern design choice, including rotary embeddings [69], QK-normalization [70], RMSNorm [71], GELU activations [72], and separate input and output embeddings. We use a context length of 512 and tokenize the dataset using the GPT-2 tokenizer. We provide additional experimental details in Appendix˜A.
@@
Why is momentum less necessary for small batch sizes? Momentum can be seen as a means of dampening oscillations that occur on ill-conditioned loss functions [73]. When we take large steps, we may overshoot the minimum across short (i.e. high curvature) axes, leading to oscillations. By averaging gradients across iterations, these oscillations cancel out in the momentum term. However, when the optimizer takes small steps, as is the case in small batch training paired with a small learning rate, parameter updates do not overshoot the minimum across short axes, so there are no oscillations to damp. We illustrate this effect in Appendix˜B through a toy example.

### v4

@@ Abstract
Figure 1: Small batch sizes are more robust to optimizer design and hyperparameter values. (a) Loss achieved by a transformer decoder-only language model with 30M active parameters trained on 600M tokens of FineWeb-Edu data using SGD, Adafactor, Adam, and Muon. At small batch sizes, all optimizers achieve similar loss; at large batch sizes, the gap between optimizers grows. (b) Loss achieved by GPT-2 (124M) trained on 2.5B tokens of FineWeb using AdamW when tuning the learning rate and $\beta_{1}$ hyperparameters for batch sizes 1 and 512. While both batch sizes achieve a similar lowest loss, the smaller batch size is much more robust to the hyperparameter values, reducing the need for hyperparameter tuning. For batch size 512, we use the default hyperparameters from gpt3 and the nanoGPT public repository [nanogpt]. For batch size 1, we turn off weight decay and rescale $\beta_{2}$ to preserve the token half-life, resulting in $\beta_{2}=0.9999$.
@@ 1 Introduction
Large batch sizes are widely believed to improve the stability of language model training [gazagnadou2019optimal, nitanda2021sharp, kim2023parameter, ghosal2023flacuna]. As a consequence, sophisticated optimizers that perform well in large-batch training are increasingly standard practice [you2017large, gupta2018shampoo, you2019large, liu2023sophia, jordanmuon, vyas2024soap]. In fact, it is common to simulate batch sizes even larger than the maximum batch size that fits into device memory through gradient accumulation [ghosal2023flacuna, kim2023parameter, zeng2023glm130b, scao2022bloom].
In small batch size pretraining experiments, one may observe loss spikes and heavy instability [xu2020optimizing, xiao2024rethinking, labonne2024finetune, wandb2025pretraining]. While it is common to decrease the learning rate for smaller batch sizes, other hyperparameters, such as the decay rates for the first and second moments in Adam ($\beta_{1}$ and $\beta_{2}$), are often held fixed across batch sizes. We show that if instead of holding $\beta_{2}$ fixed, we hold the half-life of $\beta_{2}$ fixed (measured in number of tokens), stable training is possible all the way down to batch size one, as illustrated in Figure˜1(a).
When scaling hyperparameters in this way, we find that small batch sizes can confer computational advantages and greater robustness. In the small batch regime, we observe that the speed of convergence is less sensitive to optimizer hyperparameters like the learning rate, and momentum becomes less necessary, as we find in Figure˜1(b). In contrast, large batch sizes require large learning rates to compensate for taking fewer steps at a given compute budget. Large step sizes in turn require that the optimizer make predictions about the loss surface far away from the current iterate. We find that predicting such far-away parameter updates requires a sophisticated and carefully tuned optimizer. For very small batch sizes, we find that even vanilla stochastic gradient descent (SGD), with no momentum or weight decay, becomes competitive for training language models, in contrast to findings from recent work [zhao2025deconstructing]. The ability to perform stable training without momentum can confer significant memory advantages, since we do not need to store the corresponding optimizer state, which can have a significant memory footprint. Based on our observations, we also revisit Adafactor, a memory-efficient variant of Adam, and show that it can be a compelling alternative to Adam in the small batch regime, enabling training of larger models in a memory-constrained setting.
@@
It is common in convex optimization to use higher batch sizes as the loss converges to a minimum [gazagnadou2019optimal, sievert2019improving]. Intuitively, the higher the gradient to gradient noise magnitude ratio, the smaller the batch size we should take. For example, if the gradient noise magnitude is close to zero, then we obtain roughly the same gradient estimate with a small batch size as a large batch size but pay far fewer floating-point operations (FLOPs), so small batch sizes are efficient. Smaller batch sizes allow us to take more steps for a fixed FLOP budget. As we converge to a minimum, that same ratio instead often tends to zero since the gradient norm vanishes while the gradient noise may not, in which case we may want a larger batch size. We hypothesize that language model training lives in the far-from-convergence regime where small batch sizes are efficient. Following compute-optimal scaling laws, we would not train language models into the convergence regime since we could achieve better performance by instead training a larger model on less data [kaplan2020scaling, hoffmann2022training].
@@ 2 Background: Optimization for Language Models
Stochastic gradient descent (SGD) [robbins1951stochastic, bottou2010large]. SGD updates model parameters by computing gradients on mini-batches of data as follows: $\theta_{t+1}=\theta_{t}-\eta g_{t}$, where $\theta_{t}$ are the model parameters at step $t$, $\eta$ is the learning rate, and $g_{t}=\nabla_{\theta}\mathcal{L}(\theta_{t})$ is the gradient of the loss function with respect to $\theta_{t}$ averaged over $B$ samples. We refer to $B$ as the batch size and consider the extreme case of $B=1$ in our experiments.
@@
Adam [kingma2014adam]. Adam is an adaptive optimizer that maintains an exponential moving average (EMA) of the gradient and squared gradient, referred to as the first and second moments, and denoted as $m_{t}$ and $v_{t}$. The timescales of these moving averages are controlled by decay rates $\beta_{1}$ and $\beta_{2}$:
@@
Higher values of $\beta_{1}$ and $\beta_{2}$ place more weight on past gradients, meaning that the moments are averaged over longer timescales and evolve slowly. When training language models, values like $\beta_{1}=0.9$ and $\beta_{2}\in[0.95,0.98]$ are often used in practice [gpt3, touvron2023llama]. While most researchers and practitioners do not scale $\beta_{1}$ and $\beta_{2}$ with the batch size, we show in our work that scaling $\beta_{2}$ in particular is crucial for achieving good performance with Adam at a small batch size.
@@
Adafactor [shazeer2018adafactor]. The second moment estimate $v_{t}$ used in Adam requires storing a number of floating point values equal to the parameter count of the model itself. Instead, Adafactor only stores the per-row and per-column sums of the moving average for the neural network weight matrices and estimates the per-parameter second moment using these sums. Adafactor also does not store a first moment estimate at all. Therefore, for a weight matrix with shape $d_{1}\times d_{2}$, the memory requirements for the moving averages are reduced from $2\times d_{1}\times d_{2}$ for Adam to $d_{1}+d_{2}$ for Adafactor. We demonstrate in Section˜4.5 that thanks to this sublinear memory cost, Adafactor achieves the best performance–memory trade-off.
Muon [jordanmuon]. Another optimizer that was recently demonstrated to be competitive for language model training is MomentUm Orthogonalized by Newton-Schulz (Muon). Muon is specifically designed for $\geq 2$-dimensional parameter tensors, such as weight matrices in linear layers, where the Newton-Schulz iterative method is applied to the first moment estimate $m_{t}$ before using it to update the parameters. Muon is often used in conjunction with other optimizers like Adam. In our experiments using Muon, we only apply it to hidden layers and use Adam for input embedding and final layers.
@@ 3 Related Work
Batch size has long been studied in both theory and practice: its impact on the optimization path [jastrzkebski2018relation, jastrzebski2020break, lee2022trajectory, atanasov2024optimization], its interaction with different optimizers [wilson2003general, chen2016revisiting, data_parallelism_shallue, noisy_quadratic], and its effect on the optimal learning rates and momentum hyperparameters [smith2017bayesian, granziol2022learning, wang2023marginal, fu2023and, data_parallelism_shallue, momentum_schemes, critical_batch_size_scaling, porian_resolving, critical_batch_size_scaling, palm, adam_can_converge]. Prior work has also explored related themes, such as the flatness argument favoring SGD [lecun2002efficient, keskar2016large, goyal2017accurate, masters2018revisiting], SGD’s implicit biases [arora2019implicit, barrett2020implicit, haochen2021shape, smith2021origin], and the comparison between stochastic and full-batch optimization [ge2015escaping, du2017gradient, lewkowycz2020large, geiping2021stochastic]. However, modern language model training is neither convex nor characterized by multiple epoch training like vision settings. In fact, we often train for a limited token budget and stop far before convergence, since otherwise scaling laws suggest that we could train a much better language model for the same compute budget by training a larger model on fewer tokens [hoffmann2022training]. Hence, prior intuitions, such as the need for a larger batch size in the convergence phase, may not apply. For this reason, in the rest of this section, we cover works that focus on modern language model training specifically.
@@
critical_batch_size_scaling and data_parallelism_shallue examine the critical batch size, which represents the threshold beyond which greater data parallelism would lead to diminishing returns, and tune all optimizer hyperparameters. They argue that any batch size below the critical batch size should perform equally well when training for the same number of tokens. Differently from these works, we empirically demonstrate that small batch sizes, including the extreme value of batch size one, perform on par with or even better than larger batch sizes, and we also show that small batches are more robust to optimizer and hyperparameter choices and enable SGD to perform competitively. We validate these findings on language models with modern design choices and up to 1.3 billion parameters.
@@
On the other hand, vyas2023beyond argue that gradient noise from small batch sizes does not yield similar performance advantages in the single epoch setting as it does in multiple-epoch training. Experimental results from filatov2024time also suggest that small batches do not work as well as large batches for language models trained on C4 [dodge2021documenting], but notably, the authors do not tune the decay rates $\beta_{1}$ and $\beta_{2}$ of the Adam optimizer. In the same vein, xiao2024rethinking claim that both excessively small and large batch sizes underperform for language model training with Adam, leading to a U-shaped curve of the loss as a function of the batch size. We reproduce the exact experiments performed by xiao2024rethinking and show that their conclusions are an artifact of incomplete hyperparameter tuning; we find that simply using a slower decay rate of the second moment estimate $\beta_{2}$ makes small batch sizes competitive.
@@
zhao2025deconstructing train language models on C4 using different diagonal preconditioning optimizers, namely Adam, Lion, Adafactor with momentum, SGD with momentum, and signed SGD with momentum. They demonstrate that all these optimizers, aside from SGD which lags behind, achieve comparable optimal performance and robustness to hyperparameters. We challenge this conclusion in our work and argue that in the small batch size regime, SGD can perform on par with Adam.
@@
data_parallelism_shallue, momentum_schemes and critical_batch_size_scaling find that momentum is not necessary for SGD when the batch size is small, which aligns with our findings. However, momentum_schemes focus on small-scale vision experiments, data_parallelism_shallue and momentum_schemes do not compare the performance of SGD vs. Adam, and critical_batch_size_scaling claim that SGD performs significantly worse than Adam. We hypothesize this observation could be a result of critical_batch_size_scaling using 64 as the smallest batch size – we test batch sizes all the way down to 1. noisy_quadratic highlight that the benefits of more sophisticated optimizers diminish as batch size decreases, but their theoretical results assume a convex quadratic objective and their empirical results only cover a single language modeling experiment that is limited to a two-layer transformer.
@@
porian_resolving and critical_batch_size_scaling find that increasing the value of $\beta_{2}$ is important for small-batch training with Adam. However, their approach is based on discrete hyperparameter sweeps across $\beta_{2}$. In contrast, we introduce and validate a principled scaling rule that holds the second-moment half-life constant in terms of tokens. ema_scaling propose a scaling rule for the decay coefficient of exponential moving averages similar to ours, but applied in the context of model weight averaging. adam_can_converge show that there exists a threshold of $\beta_{2}^{*}$ such that when $\beta_{2}$ > $\beta_{2}^{*}$, Adam converges (this threshold decreases with batch size, i.e. the threshold is higher for smaller batch sizes). palm note that increasing $\beta_{2}$ during training leads to better performance for rare token embeddings. While this is an important insight, we provide a simpler batch-size-aware heuristic that applies even when using a constant $\beta_{2}$ throughout training.
@@ 4 Experimental Results
Recent works find that larger batch sizes and relatively sophisticated optimizers are necessary for stable language model training. For instance, zhao2025deconstructing, adam_sgd_lm, and transformers_need_adam observe that SGD has an extremely slow convergence speed for training language models in comparison to adaptive optimizers like Adam. Other works find that Adam performs poorly with small batch sizes [vyas2023beyond, xiao2024rethinking, liu2024epochal, li2024surge].
In this section, we revisit and challenge these beliefs by noting that such findings arise due to poor hyperparameter choices and specific experimental setups, for example, not adjusting the decay parameters $\beta_{1}$ and $\beta_{2}$ for Adam at small batch sizes, or using too large a batch size for SGD. We put these beliefs to the test by running a thorough grid search over optimizer hyperparameters and batch sizes for SGD, Adam, Adafactor, and Muon. As the cost of a grid search grows exponentially with the number of hyperparameters, we only perform exhaustive grid searches on a small transformer decoder-only model with 30 million active parameters following nanodo and xiao2024rethinking. We use these results to propose heuristics for scaling hyperparameters across batch sizes, and we validate our findings at a larger scale by pretraining GPT-2 (124M) [gpt2] and GPT-3 (1.3B) [gpt3] and fine-tuning Gemma 3 (4B) [gemma3]. In our experiments, we study batch sizes spanning more than four orders of magnitude: $\{1,4,16,64,256,1024,\text{ and }4096\}$.
@@ 4.1 Small Batch Sizes Render Sophisticated Optimizers Unnecessary
We train the 30M model on 600 million tokens from the FineWeb-Edu dataset [penedo2024fineweb], following Chinchilla scaling laws [hoffmann2022training]. Our model adopts modern design choice, including rotary embeddings [rope], QK-normalization [qknorm], RMSNorm [rmsnorm], GELU activations [gelu], and separate input and output embeddings. We use a context length of 512 and tokenize the dataset using the GPT-2 tokenizer. We provide additional experimental details in Appendix˜A.
@@
Why is momentum less necessary for small batch sizes? Momentum can be seen as a means of dampening oscillations that occur on ill-conditioned loss functions [goh2017why]. When we take large steps, we may overshoot the minimum across short (i.e. high curvature) axes, leading to oscillations. By averaging gradients across iterations, these oscillations cancel out in the momentum term. However, when the optimizer takes small steps, as is the case in small batch training paired with a small learning rate, parameter updates do not overshoot the minimum across short axes, so there are no oscillations to damp. We illustrate this effect in Appendix˜B through a toy example.